# Benchmarking the Spatial Robustness of DNNs via Natural and Adversarial Localized Corruptions

### v1

@@ 4 Localized Natural Corruptions
Figure 1: Illustrations of the proposed framework for evaluating localized natural corruptions. A configuration is specified to define the settings for the localized corruption analysis, and a single sample from the dataset is returned as a tuple consisting of the original image, the original ground truth label, and the corrupted image, along with the corrupted region that highlights which areas have been perturbed.
This section presents the framework we designed to evaluate localized natural corruptions. Our framework was inspired by ImageNet-C [14], which however does not support localized corruptions. We consider different transformations $c$, which can be applied at various severity levels $s$ (i.e., 1, 2, 3, 4, 5), as selected by the user. A natural corruption applied to an input $x$ with severity level $s$ is denoted by $c(x)=\Gamma(x,s)$, where $\Gamma$ represents a natural transformation, such as Gaussian noise, blur, motion blur, etc.
The corrupted region $M$ is automatically selected according to the following pipeline. The image $x$ is first divided into non-overlapping patches with size $\Delta=(\Delta_{x},\Delta_{y})$, thereby obtaining a total of $P=P_{x}\times P_{y}$ patches with $P_{x}=\lceil W/\Delta_{x}\rceil$, $P_{y}=\lceil H/\Delta_{y}\rceil$, where $H$ and $W$ are the height and width of $x$, respectively. Given a specified corruption ratio $r$, each patch, and therefore all its pixels, is selected to be perturbed with probability $r$. All pixels within a selected patch are assigned a value of $1$ in $M$, while the others have a value of 00.
In Figure 1, we present an illustration of a possible configuration used within the tested framework and the output of a dataset sample in the corrupted validation set of Cityscapes [8]. Specifically, a sample corresponds to a tuple containing the original image and its label (kept available to allow comparisons during the evaluation), along with the perturbed image and the corruption region mask $M$. During evaluation, for each image, the set of patches to be perturbed (and thus the mask $M$) is recomputed, effectively randomizing the selection of the perturbed areas for the evaluation.
@@ 5 Localized Adversarial Corruptions
This section presents our novel analysis for localized adversarial perturbations. Although previous work already addressed the study of localized adversarial perturbations against SS models [24], we next highlight the difficulties of performing a comprehensive spatial adversarial evaluation using a single attack. To this end, we begin by discussing classic approaches and pointing out their drawbacks. Finally, we introduce our region-aware multi-attack algorithm designed to address these issues and offer a better understanding of the worst-case spatial robustness of ML models.
@@ 5.1 Attacks for SS from previous work
In the context of classic adversarial perturbations [32, 3], the main idea is to craft a specific adversarial noise $\delta$ that realizes a worst-case perturbation constrained by a magnitude $\varepsilon$ under a specific norm, e.g., the $l_{\infty}$ norm, such that $\|\delta\|_{\infty}\leq\varepsilon$. Adversarial perturbations are typically computed by optimizing an adversarial loss function $\mathcal{L}$, e.g., cross-entropy in image classification, which aims to find a perturbation $\delta$ for an input $x$ that forces the model to make incorrect predictions $\neq y$:
@@
After recalling Equation 1, it is possible to express inputs affected by localized attacks with a perturbation $\delta$ and a mask $M$ as follows:
@@
When interested in dense prediction tasks, such as semantic segmentation, where predictions are made at the pixel level, the perturbation is generally computed to maximize the misclassification of all or a subset of pixels in the image [24, 29]. In the latter case, it is possible to define a fooling region as a subset of the input space for which we intend to induce a misprediction. The fooling region can be arbitrarily defined by a mask $F\in\{0,1\}^{H\times W}$ that does not necessarily need to overlap with the corruption mask $M$. In this context, the perturbation $\delta$ can be optimized as follows:
@@
where $\mathcal{L}_{\mathcal{F}}$ is an adversarial loss designed to focus the model performance in the fooling region $F$ only. In practice, to solve the above optimization problem, the perturbation $\delta$ can be iteratively updated with step size $\alpha$ to maximize the classification error:
@@ 5.2 Limitations of previous work
Although previous approaches are capable of crafting strong perturbations by corrupting as many pixels as possible, we argue that comprehensively evaluating the spatial robustness of semantic segmentation models requires multiple instances of attacks applied to the same area. A single perturbation, even when computed using a strong attack, cannot cover all regions susceptible to misprediction. This happens because, as shown in Eq. 9, crafting an optimal localized adversarial perturbation involves solving a complex multi-objective optimization problem, where the goal is to increase the loss for each pixel in the image, representing a wide range of distinct objectives. Many of these objectives are difficult to solve simultaneously, often leading to suboptimal or locally optimal solutions. To support this observation, we provide representative analysis results in Figures 2(a) and  2(b) (discussed in the next paragraphs), which highlight, from both inter-class and intra-class perspectives, that adversarial attacks targeting specific subsets of pixels may not generalize effectively to others, making it particularly challenging to craft a single localized perturbation that generates worst-case adversarial effects.
@@
Figure 2: Analysis of limitations in SS adversarial optimization for single localized perturbations: (a) Inter-class analysis: Attacked images and predictions across different target classes. (b) Intra-class analysis: Targeted multi-attack on the class ‘Car’. Visualization of adversarial effects in different areas. Each subfigure highlights a distinct aspect of the attack.
@@
Let us first focus on inter-class attacks with reference to Figure 2(a), where localized perturbations are applied to the center of the image ($100\times 100$, using iterative targeted attacks [24] with $\epsilon=32/255$). In this example, attacks are performed each targeting different classes in separate runs (’road’, ’sky’, and ’car’). The outputs of the addressed targeted attacks and their corresponding classes are shown from the third to fifth subfigure. Additionally, we also show the output prediction of an untargeted attack (targeting all classes), in the last subfigure. The results demonstrate that the adversarial patterns produce distinct adversarial effects, highlighting that the multi-objective optimization landscape for different target classes is highly divergent, evidencing difficulties for a targeted attack to affect the areas interested by others. Furthermore, when observing the output of the untargeted attack, it closely resembles the output of the ’road’ attack, suggesting a bias in the multi-objective optimization process, possibly given by the fact that the majority of pixels of the image are classified as ’road’.
@@
We then analyze the same problem in the context of intra-class attacks. In this case, as shown in Figure 2(b), we crafted attacks to misclassify pixels belonging to the ’car’ class. We started with an initial fooling region, highlted in the first subfigure, consisting of all pixels labeled as ’car’ in the ground truth. After executing the attack of Eq. (9) a first time (denoted by attack 1) using the initial fooling region, we were capable of misclassifying a limited subset of the targeted pixels only (note the attack output in the second row and the corresponding red pixels in the third row of Figure 2(b)). To carry out a second attack (attack 2), we excluded from the initial fooling area the pixels that were successfully attacked in the previous attack. This adjustment allows us to investigate whether previously attacked pixels negatively influence the optimization process for the remaining pixels classified as ’car’. The restricted fooling region for the second attack is shown in the third row. As it can be observed from Figure 2(b), the second attack expands the attacked area, indicating that previously attacked regions can indeed hinder the optimization process when exploring new attackable regions. This process is repeated for a third attack (last columns in the second and third rows of the figure), confirming the previous observation.
@@
Finally, for completeness, we present the cumulative adversarial output in the third subfigure (first row, third column). This cumulative output assigns each pixel its adversarial misclassified prediction if at least one attempted attack has misclassified it (the first attack is considered for simplicity); otherwise, it retains the original model prediction. This approach provides a clearer worst-case understanding of whether each given pixel can be misclassified by at least one attack and will be used in the algorithm proposed in the next subsection.
@@ 5.3 Achieving Stronger Spatial Adversarial Analysis
The observations made in the previous section allow concluding that single instances of SS attacks provide an incomplete assessment of the worst-case spatial adversarial robustness of a model. To address this limitation and establish a more comprehensive benchmark for evaluating SS models under localized adversarial perturbations, we propose a new method, called region-aware multi-attack, in Algorithm 1. This algorithm builds upon the observations made above and is next explained step-by-step. At a high level, the method iteratively attacks the input while progressively modifying the attacked area by refining the fooling region at each attack run, thereby maximizing the coverage of attacked pixels.
Algorithm 1 Region-aware Multi-Attack Analysis
@@
3:$\hat{y}_{\text{clean}}\leftarrow\text{argmax}(f(x))$triangleright Clean preds
4:$\hat{y}_{\text{O}}\leftarrow\hat{y}_{\text{clean}}$triangleright Initialize cumulative output
5:$M_{\text{corr}}\leftarrow(\hat{y}_{\text{O}}=\textbf{y})$triangleright Mask of correct preds
6:$\mathcal{F}\leftarrow M_{\text{corr}}$triangleright Initial fooling region
@@
8:     $\mathcal{F}\leftarrow\mathcal{F}\cdot M_{\text{corr}}$triangleright Update fooling region
9:     $x_{\text{adv}}\leftarrow\textit{Attack}(x,\textbf{y},f,\mathcal{F})$
10:     $\hat{y}_{\text{adv}}\leftarrow\text{argmax}(f(x_{\text{adv}}))$triangleright Preds after attack
11:     $M_{\text{cur}}\leftarrow(\hat{y}_{\text{adv}}=\textbf{y})$triangleright Current correct pixels
12:     $\bar{M}\leftarrow(1-M_{\text{cur}})\cdot M_{\text{corr}}$triangleright Newly misclassified pixels
13:     $M_{\text{corr}}\leftarrow M_{\text{cur}}\cdot M_{\text{corr}}$triangleright Update correct preds mask
14:     $\hat{y}_{\text{O}}\leftarrow M_{\text{corr}}\cdot\hat{y}_{\text{O}}+\bar{M}\cdot\hat{y}_{\text{adv}}$triangleright Update cumul. output
@@
16:return $\textit{Eval}(\hat{y}_{\text{O}},\textbf{y})$
The proposed algorithm takes as input a given image $x$, its labels y, and an SS model $f$, and aims to evaluate the robustness of the image by running multiple instances of attacks ($N_{\textit{Att}}$) based on a given attack method (referred to as Attack in the algorithm). At the beginning of the algorithm, the cumulative output is initialized to the correct model predictions, and the initial fooling region consists of all pixels that are correctly predicted in the non-perturbed output (line 4) [^1] We address the untargeted case for simplicity and to generalize as much as possible, covering as many pixels as possible. However, the approach can be easily extended to a targeted formulation.. Subsequently, the attack iteratively targets the pixels in the fooling region using a classic iterative attack (e.g., Equation 10, line 7), computes the corresponding output (line 8), and updates the fooling region by identifying pixels in the preceding fooling region that have not yet been misclassified in the current attack (lines 9–12 and line 6). At the end of each attack, the cumulative adversarial output is updated to replace the predictions of newly misclassified pixels that had previously been correctly classified in earlier attacks.
@@
Please note that the proposed formulation is entirely untargeted. As demonstrated in the experimental section, the approach also addresses class-wise coverage by incorporating an incremental fooling region. This strategy helps mitigate intra-class issues that may arise even in targeted attacks, as discussed above and illustrated in Figure 2(b).
@@ 6.1 Experimental Setup
The experiments were conducted with NVIDIA A100 GPUs, using the validation set of Cityscapes [9], which serves as the reference driving dataset for high-resolution scene-understanding segmentation. In the context of semantic segmentation, this dataset has frequently been used for robustness evaluation, as it addresses real-world and complex outdoor scenarios [29, 24, 28]. The Cityscapes validation set includes 500 high-resolution images with an original resolution of ($1024\times 2048$), resized to manage computational costs (as done in [24]). We used this set to evaluate both localized natural corruptions and adversarial attacks.
Models. We targeted multiple models for semantic segmentation, all known for their ability to achieve high performance on large images while maintaining practical inference times affordable for many real-time applications. Specifically, we considered ICNet [39], BiseNet [36], DDRNet [15], PSPNet [40], SegFormer [34], PIDNet [35], and DeepLabV3 [7], including multiple versions of each model to evaluate robustness differences, such as variations in the backbone architecture. The average accuracy of these models on the clean validation set of Cityscapes is reported in the first column of Table 1.
Metrics. The spatial robustness analysis was conducted by evaluating the pixel-wise accuracy of the models in the con-corrupted region $A_{M}$ and the corrupted region $A_{\overline{M}}$, both defined in Section 3. Furthermore, to better understand the robustness degradation as the evaluation parameters vary, we analyzed the previous scores in terms of the relative corruption error (RCE) with respect to the accuracy of the addressed models in the same area of interest when no natural corruption or attacks are applied. Leveraging the localized corruption accuracy defined in Eq. (4), in the case of the relative corruption error in the corrupted region the relative error can be computed as
$\text{RCE}(f,C,M)=\frac{A_{\mathcal{P}}(f,\{\emptyset\},M)-A_{\mathcal{P}}(f,C,M)}{A_{\mathcal{P}}(f,\{\emptyset\},M)},$
were $\mathcal{P}_{i}(M)=M_{i}/|M|$. A similar definition can be derived for the RCE outside the corrupted region. Note also that, as indicated in Section 3, natural corruptions have been evaluated across different regions (see Equation (6)).

### v3

@@ 4 Localized Natural Corruptions
Figure 1: Illustrations of the proposed framework for evaluating localized natural corruptions. On the left, two example configurations define the settings for localized corruption analysis: patch size, corruption ratio, corruption type, and the severity level. The framework returns a dataset sample as a tuple consisting of the original image, the ground truth label, the corrupted image, and a mask highlighting the corrupted areas.
This section presents the framework designed to evaluate localized natural corruptions. The framework includes all the corruption types available in ImageNet-C [hendrycks2018benchmarking] which, however, does not support localized corruptions. In particular, we consider different transformations $c$, which can be applied at various severity levels $s$ (i.e., 1, 2, 3, 4, 5), as selected by the user. A natural corruption applied to an input $x$ with severity level $s$ is denoted by $c(x)=\Gamma(x,s)$, where $\Gamma$ represents a natural transformation, such as gaussian noise, gaussian blur, motion blur, etc.
To enable a systematic, reproducible, and broad evaluation of spatial robustness, the corrupted region indicated by $M$ is automatically selected according to the following pipeline. The image $x$ is first divided into non-overlapping patches with size $\Delta=(\Delta_{x},\Delta_{y})$, thereby obtaining a total of $P=P_{x}\times P_{y}$ patches with $P_{x}=\lceil W/\Delta_{x}\rceil$, $P_{y}=\lceil H/\Delta_{y}\rceil$, where $H$ and $W$ are the height and width of $x$, respectively. Given a specified corruption ratio $r$, each patch, and therefore all its pixels, is selected to be perturbed with probability $r$. All pixels within a selected patch are assigned a value of $1$ in $M$, while the others have a value of 0.
In Figure 1, we present an illustration of a possible configuration used within the tested framework and the output of a dataset sample in the corrupted validation set of Cityscapes [cordts2016cityscapes]. Specifically, a sample corresponds to a tuple containing the original image and its label (kept available to allow comparisons during the evaluation), along with the perturbed image and the corruption region mask $M$. During evaluation, for each image, the set of patches to be perturbed (and thus the mask $M$) is recomputed, effectively randomizing the selection of the perturbed areas for the evaluation.
@@ 5 Localized Adversarial Corruptions
This section presents our novel analysis for localized adversarial perturbations. Although previous work already addressed the study of localized adversarial perturbations against SS models [nakka_indirect_2020], we next highlight the difficulties of performing a comprehensive spatial adversarial evaluation using a single attack. To this end, we begin by discussing classic approaches and pointing out their drawbacks. Finally, we introduce our region-aware multi-attack algorithm designed to address these issues and offer a better understanding of the worst-case spatial robustness of computer vision models.
@@ 5.1 Attacks for SS from previous work
In the context of classic adversarial perturbations [Szegedy14, biggio2018wild], the main idea is to craft a specific adversarial noise $\delta$ that realizes a worst-case perturbation constrained by a magnitude $\varepsilon$ under a specific norm, e.g., the $l_{\infty}$ norm, such that $\|\delta\|_{\infty}\leq\varepsilon$. Adversarial perturbations are typically computed by optimizing an adversarial loss function $\mathcal{L}$, e.g., cross-entropy in image classification, which aims to find a perturbation $\delta$ for an input $x$ that forces the model to make incorrect predictions:
@@
Recalling Equation 1, it is possible to express inputs affected by localized attacks with a perturbation $\delta$ and a mask $M$ as follows:
@@
When interested in dense prediction tasks, such as semantic segmentation, where predictions are made at the pixel level, the perturbation is generally computed to maximize the misclassification of all or a subset of pixels in the image [nakka_indirect_2020, rossolini_tnnls_2023]. In the latter case, it is possible to define a fooling region as a subset of the input space for which we intend to induce a misprediction. The fooling region can be arbitrarily defined by a mask $F\in\{0,1\}^{H\times W}$ that does not necessarily need to overlap with the corruption mask $M$. In this context, the perturbation $\delta$ can be optimized as follows:
@@
where $\mathcal{L}_{\mathcal{F}}$ is an adversarial loss designed to focus the model performance in the fooling region $F$ only. In practice, to solve the above optimization problem, the perturbation $\delta$ can be iteratively updated with step size $\alpha$ to maximize the classification error [nakka_indirect_2020, pgd_attack]:
@@ 5.2 Limitations of previous work
Although previous approaches are capable of crafting strong perturbations by corrupting as many pixels as possible, we argue that comprehensively evaluating the spatial robustness of semantic segmentation models requires multiple attacks applied to the same area. A single perturbation, even when computed using a strong attack, cannot cover all regions susceptible to misprediction. This happens because, as shown in Equation 9, crafting an optimal localized adversarial perturbation involves solving a complex multi-objective optimization problem, where the goal is to increase the loss for each pixel in the image, representing a wide range of distinct objectives. Many of these objectives are difficult to solve simultaneously, often leading to suboptimal or locally optimal solutions. To support this observation, we provide representative analysis results in Figures 2(a) and  2(b) (discussed in the next paragraphs), which highlight, from both inter-class and intra-class perspectives, that adversarial attacks targeting specific subsets of pixels may not generalize effectively to others, making it particularly challenging to craft a single localized perturbation that generates worst-case adversarial effects.
@@
Figure 2: Analysis of limitations for single localized perturbations in $S$: (a) Inter-class analysis: clean prediction, attacked image, and predictions across different target classes, shown in sequence. (b) Intra-class analysis: Targeted multi-attack on the class ‘Car’. The first row displays the fooling area, the attacked image, and the cumulative adversarial output. The second row shows the output after each attack iteration, while the last row visualizes how the fooling region is selected and which pixels are actually affected at each attack.
@@
Let us first focus on inter-class attacks with reference to Figure 2(a), where localized perturbations are applied to the center of the image ($100\times 100$, using iterative targeted attacks [nakka_indirect_2020] with $\epsilon=32/255$). In this example, attacks are performed each targeting different classes in separate runs (‘road’, ‘sky’, and ‘car’). The outputs of the addressed targeted attacks and their corresponding classes are shown from the third to the fifth subfigures. Additionally, we also show the output prediction of an untargeted attack (targeting all classes), in the last subfigure. The results demonstrate that the adversarial patterns produce distinct adversarial effects. This highlights that the multi-objective optimization landscape for different target classes is highly divergent, evidencing difficulties for a targeted attack to affect the areas interested by others. Furthermore, when observing the output of the untargeted attack, it closely resembles the output of the ‘road’ attack, suggesting a bias in the multi-objective optimization process, possibly given by the fact that the majority of pixels of the image are classified as ‘road’.
@@
We then analyze the same problem in the context of intra-class attacks. In this case, as shown in Figure 2(b), we crafted attacks to misclassify pixels belonging to the ‘car’ class. We started with an initial fooling region, highlighted in the first subfigure, consisting of all pixels labeled as ‘car’ in the ground truth. After executing the attack of Equation 9 a first time (denoted by ‘attack 1’) using the initial fooling region, we were capable of misclassifying a limited subset of the targeted pixels only (note the attack output in the second row and the corresponding red pixels in the third row of Figure 2(b)). To carry out a second attack (attack 2), we excluded from the initial fooling area the pixels that were successfully attacked in the previous attack. This adjustment allows us to investigate whether previously attacked pixels negatively influence the optimization process for the remaining pixels classified as ‘car’. The restricted fooling region for the second attack is shown in the third row. As can be observed from Figure 2(b), the second attack expands the attacked area, indicating that previously attacked regions can indeed hinder the optimization process when exploring new attackable regions. This process is repeated for a third attack (last columns in the second and third rows of the figure), confirming the previous observation.
@@
Finally, for completeness, we present the cumulative adversarial output in the third subfigure (first row, third column). This cumulative output assigns each pixel its adversarial misclassified prediction if at least one attempted attack has misclassified it (the first misclassification is considered for simplicity); otherwise, it retains the original model prediction. This approach, provides a clearer worst-case understanding of whether each given pixel can be misclassified by at least one attack and will be used in the algorithm proposed in the next subsection.
@@ 5.3 Achieving Stronger Spatial Adversarial Analysis
The observations made in the previous section allow concluding that classically applied single instances of SS attacks provide an incomplete assessment of the worst-case spatial adversarial robustness of a model. To address this limitation and assess a more comprehensive benchmark under localized adversarial perturbations, a method called region-aware multi-attack adversarial analysis is proposed in Algorithm 1. The algorithm builds upon the observations made above and is next explained step-by-step. At a high level, the method iteratively generates a new localized adversarial attack on a region $M$ while progressively refining the fooling region $\mathcal{F}$ at each attack run, thereby maximizing the coverage of attacked pixels.
Algorithm 1 Region-aware Multi-Attack Adversarial Analysis
@@
[…]
@@
8:  $\mathcal{F}\leftarrow\mathcal{F}\cdot\Psi_{\text{corr}}$ $\triangleright$ Update fooling region
9:  $x_{\text{adv}}\leftarrow\textit{Attack}_{M}(x,y,f,\mathcal{F})$
10:  $\hat{y}_{\text{adv}}\leftarrow\text{argmax}(f(x_{\text{adv}}))$ $\triangleright$ Preds after attack
11:  $\Psi_{\text{cur}}\leftarrow(\hat{y}_{\text{adv}}=y)$ $\triangleright$ Current correct pixels
12:  $\bar{\Psi}\leftarrow(1-\Psi_{\text{cur}})\cdot\Psi_{\text{corr}}$ $\triangleright$ Newly misclassified pixels
13:  $\Psi_{\text{corr}}\leftarrow\Psi_{\text{cur}}\cdot\Psi_{\text{corr}}$ $\triangleright$ Update correct aux. mask
14:  $\hat{y}_{\text{O}}\leftarrow\Psi_{\text{corr}}\cdot\hat{y}_{\text{O}}+\bar{\Psi}\cdot\hat{y}_{\text{adv}}$ $\triangleright$ Update cumulative output
@@
16:return $\textit{Eval}(\hat{y}_{\text{O}},y)$
The proposed algorithm takes as input a given image $x$, its labels $y$, and an SS model $f$, and aims to evaluate the robustness of the image by running multiple attacks ($N_{\textit{Att}}$), applied to the same area $M$, based on a given method (referred to as $\textit{Attack}_{M}$ in the algorithm). At the beginning of the algorithm, the cumulative output is initialized to the correct model predictions, and the initial fooling region consists of all pixels that are correctly predicted in the non-perturbed output (line 4) [^2] We address the untargeted case for simplicity and to generalize as much as possible, covering as many pixels as possible. However, the approach can be easily extended to a targeted formulation.. Subsequently, the attack iteratively targets the pixels in the fooling region using a classic iterative attack (e.g., Equation 10, line 7), computes the corresponding output (line 8), and updates the fooling region by identifying pixels in the preceding fooling region that have not yet been misclassified in the current attack (lines 9–12 and line 6). At the end of each attack, the cumulative adversarial output is updated to replace the predictions of newly misclassified pixels that had previously been correctly classified in earlier attacks.
@@
Please note that, while the proposed formulation is entirely untargeted, this strategy also helps mitigate intra-class issues that may arise even in targeted attacks, as discussed above and illustrated in Figure 2(b).
@@ 6.1 Experimental Setup
The experiments were conducted with NVIDIA A100 GPUs, using the validation set of Cityscapes [DBLP:conf/cvpr/CordtsORREBFRS16], which serves as the reference driving dataset for high-resolution scene-understanding segmentation. In the context of semantic segmentation, this dataset has frequently been used for robustness evaluation, as it addresses real-world and complex outdoor scenarios [rossolini_tnnls_2023, nakka_indirect_2020, rossolini_iccps24]. The Cityscapes validation set includes 500 high-resolution images with an original resolution of ($1024\times 2048$), resized to manage computational costs (as done in [nakka_indirect_2020]). We used this set to evaluate both localized natural corruptions and adversarial attacks.
Models. We targeted multiple models for semantic segmentation, all known for their ability to achieve high performance on large images while maintaining practical inference times affordable for many real-time applications. Specifically, we considered ICNet [icnet_paper], BiseNet [bisenet_paper], DDRNet [ddrnet_paper], PSPNet [pspnet_paper], SegFormer [xie2021segformer], PIDNet [xu2023pidnet], and DeepLabV3 [chen2017rethinking_deeplab], including multiple versions of each model to evaluate robustness differences, such as variations in the backbone architecture. Pretrained versions, available in the paper’s repository, were used for the experimental analysis. The average accuracy of these models on the clean validation set of Cityscapes is reported in the first column of Table 1. For fairness, all results reported for evaluating natural robustness are averaged over 5 runs using the same set of seeds across model comparisons.
Metrics. The spatial robustness analysis was conducted by evaluating the pixel-wise accuracy of the models in the non-corrupted region $A_{\overline{M}}$ and the corrupted region $A_{M}$, both defined in Section 3. Furthermore, to better understand the robustness degradation as the evaluation parameters vary, we analyzed the previous scores in terms of the relative corruption error (RCE) with respect to the accuracy of the addressed models in the same area of interest when no natural corruption or attacks are applied. Leveraging the localized corruption accuracy defined in Equation 4, the relative corruption error can be computed as
$\text{RCE}(f,C,M)=\frac{A_{\mathcal{P}}(f,\{\emptyset\},M)-A_{\mathcal{P}}(f,C,M)}{A_{\mathcal{P}}(f,\{\emptyset\},M)},$ (11)
where $\mathcal{P}_{i}(M)=(1-M_{i})/\overline{\mathcal{S}}$. A similar definition can be derived for the RCE within the corrupted region, considering $\mathcal{P}_{i}(M)=M_{i}/{\mathcal{S}}$. Note also that, as indicated in Section 3, natural corruptions have been evaluated across different regions (see Equation 6).
6.2 Evaluation of Natural Robustness